Find the solution of (3x-y+6)dx +(6x-2y-6)dy=0 using Case 1

1
Expert's answer
2021-12-20T13:36:28-0500
"(3x-y+6)dx +(6x-2y-6)dy=0"

"\dfrac{dy}{dx}=-\dfrac{(3x-y)+6}{2(3x-y)-6}"

Let "v=3x-y." Then


"\dfrac{dv}{dx}=3-\dfrac{dy}{dx}=3+\dfrac{v+6}{2v-6}=\dfrac{7v-12}{2v-6}"

"\dfrac{2v-6}{7v-12}dv=dx"

Integrate


"\int\dfrac{2v-6}{7v-12}dv=\int dx"

"\int\dfrac{2v-6}{7v-12}dv=\dfrac{2}{7}\int\dfrac{7v-12}{7v-12}dv-\dfrac{18}{7}\int\dfrac{1}{7v-12}dv"

"=\dfrac{2}{7}v-\dfrac{18}{49}\ln(|7v-12|)+C_1"

"\dfrac{2}{7}v-\dfrac{18}{49}\ln(|7v-12|)=x-\dfrac{C}{49}"

"\dfrac{2}{7}(3x-y)-\dfrac{18}{49}\ln(|21x-7y-12|)=x-\dfrac{C}{49}"

"49x-42x+14y+18\ln(|21x-7y-12|)=C"

The solution is


"7x+14y+18\ln(|21x-7y-12|)=C"